Assume that the mean hourly cost to operate a commercial airplane

follows the normal distribution with a mean of $2,100 per hour and a standard deviation of $250.

What is the operating cost for the lowest 3 percent of the airplanes?


1
Expert's answer
2021-07-13T09:50:19-0400

Let "X=" the operating cost: "X\sim N(\mu, \sigma^2)."

Given "\mu=2100, \sigma=250."

"P(X<\dfrac{x-2100}{250})=0.03"

"\dfrac{x-2100}{250}=-1.8808"

"x=250(-1.8808)+2100"

"x=\$1630"

The operating cost for the lowest 3 percent of the airplanes is $1630.